OpenStudy (anonymous):

F(x)= 7xln(x) find y' and y"

OpenStudy (agent0smith):

Remember the product rule? Think of the first function as the 7x, and the other function as lnx\[\Large y = (7 x) \ln(x)\] Derivative of 7x is 7, derivative of lnx is 1/x... from the product rule: \[\Large y' = 7 \ln x + (7x) \frac{ 1 }{ x }\]

OpenStudy (anonymous):

yeah, so it would be the same thing for quotient rule right. I had a lecture on finding the derivatives of logs but i did not understand them

OpenStudy (anonymous):

is the second derivative 7/x

OpenStudy (agent0smith):

Derivative of a ln function is \[\Large \frac{ d }{ dx } \ln f(x) = \frac{ f \prime (x) }{ f(x) }\] or the derivative of the inside function, divided by the inside function.

OpenStudy (anonymous):

oh ok its the same thing as 1/f(x) * f'(x) right

OpenStudy (agent0smith):

Correct :) for y'', you're right that it's 7/x. Since \[\Large y' = 7 \ln x + (7x) \frac{ 1 }{ x } = 7 \ln x + 7\] \[\large y \prime = 7 \ln x + 7\] derivative of the 7 is just zero, so \[\Large y \prime \prime = 7 \left( \frac{ 1 }{ x }\right)\]

OpenStudy (anonymous):

ok thanks :)